OpenStudy (anonymous):

17) Solve using the elimination method 6x-9y=13.5 7y-3x=-8

OpenStudy (anonymous):

multiply second equn by 2 14y-6x = -16 6x-9y = 13.5 adding these eliminates x: 5y = -2.5 y = -0.5 plug this in second equation: 6x + 4.5 = 13.5 6x = 9 x = 1.5 (x.y) = (1.5, - 0.5)